Answer questions 1 and 10 based on a global economy of two Ricardian economies (Home, Foreign) whose industry technologies are given in the following table Product Home Foreign aLX = 0.5 0.5 Table 1: Labor hours per unit of outputs Each of these two economies is endowed with 500 labor hours (i.e. L- L*-500) 8. Again, suppose that both Home and Foreign move from autarky to a free-trade regime and that the trade price of product X is 0.4Y. If Home exports 100 units of Y to Foreign, then from this bilateral trade, Homes and Foreigns gains worth product X, respectively and units of A) 50; 75 B) 75; 100 C) 50; 125 D) 75; 150 9. Again, suppose that both Home and Foreign move from autarky to a free-trade regime and that the trade price of product X is 0.4Y. If Home exports 100 units of Y to Foreign, then from this bilateral trade, Homes and Foreigns gains worth product Y, respectively and units of A) 10; 15 B) 15; 20.5 C) 20; 37.5 D) 25; 40 10. Again, suppose that both Home and Foreign move from autarky to a free-trade regime and that the trade price of product X is 0.5Y. If Home exports 100 units of Y to Foreign, then from this bilateral trade, Homes and Foreigns gains worth product X, respectively and units of A) zero; 100 B) 100; zero C) zero; 50 D) 50; zer<o
Answer #1

8) The answer should be gain for home is 50 X and gain for foreign is 150X. Initially Home can produce 1000 units of X or 500 units of Y. Similarly foreign can produce 1000 units of X and 250 units of Y. In this case in home country 1 unit of Y equivalent to 2 units of X. In foreign country 1 unit of Y 4 units of X. If home wants to produce 100 units of Y then home has to sacrifice 200 units of X. So under auterky situation home can get 200 units of X by giving 100 units of Y. Similarly in foreign if 100 units Y to produce foreign has to sacrifice 400 units of X.

Now under trade 1X =0.4Y . It means 100 Y =250X. So now if home exports 100 units of home will get 250 units of X. Initially Home was getting 200 units of X in auterky situation. So gain for home is 250-200=50 units of X. In foreign they are getting 100 units Y just giving 250 units of X. The foreign was sacrificing 400 units of X to produce 100 units of Y. So now foreign is gaining 400-250 =150 units of X. So answer should be gain for home is 50 units of X and gain for foreign is 150 units of X.

9) The answer will be (a) . Because under auterky situation there was 1000X =500Y for home because by 500 units of labour home can produce 1000 units of X (500/0.5) or 500 units of Y. So we can say 1X =0.5 Y under auterky, now under trade 1X = 0.4Y. So gain for home in terms of Y is (0.5-0.4)*100=10 Y. Now in foreign under auterky situation 1Y =4X or 1 X =0.25 Y. So gain for foreign is (0.4-0.25)100 =0.15*100=15. So gain for home is 10 Y and for foreign is 15 Y.

10) The answer is C. Because in home country 1X =0.5 Y iniauterky situation andain trade situation it is also same. So home country will not gain from here. In the foreign 1X =0.25 Y under auterky situation. If foreign wants to produce 100 units of foreign has to sacrifice 250 units of X. But under trade foreign is getting 100 units of Y by giving 200 units of X. Because under trade 1X =0.5 Y , therefore 100 Y =200 X. So foreign gain is 250-200 =50 units of X. So gain for home is zero and gain for foreign is 50 units of X.